OpenStudy (vane11):

I'd like to know the process to get to the answer please (included) The total weekly cost (in dollars) incurred by Lincoln Records in pressing x compact discs is given by the following function. C(x) = 2000 + 2x - 0.0001x2 (0 < (or = to) x < (or = to) 6000) (a) What is the actual cost incurred in producing the 971st and the 1971st disc? (Round to the nearest cent.) 971st disc $1.81 1971st disc $1.61 (b) What is the marginal cost when x = 970 and 1970? (Round to the nearest cent.) 970 $1.81 1970 $1.61

OpenStudy (anonymous):

don't you plug in the number?

OpenStudy (anonymous):

\[C(x) = 2000 + 2x - 0.0001x^2\]\[C(971)=2000+2(971)-.0001(971)^2\]?

OpenStudy (vane11):

Did you try it on your calculator?? I tried it and I'm way off so I might be putting it in wrong if you got it like that

jimthompson5910 (jim_thompson5910):

part a looks like you're just substituting like satellite73 is showing part b will involve the derivative of C(x)

OpenStudy (anonymous):

or just the slope

OpenStudy (anonymous):

oh yeah i am way way off

OpenStudy (anonymous):

perhaps it is \[C(971)-C(970)\] for the actual cost of that disc lets try it

OpenStudy (anonymous):

C is the total cost up until that disc so the cost of producing that disc is the to total cost up to that one minus the total cost up until the one beofore

OpenStudy (vane11):

Ohhhh ok yay! haha

OpenStudy (anonymous):

i believe this is also the marginal cost, unless you are using calculus

OpenStudy (vane11):

Well the answers are the same for A and B so I'm guessing so, thanks!

OpenStudy (anonymous):

k they might be the same in any case due to rounding

OpenStudy (anonymous):

\[C(x) = 2000 + 2x - 0.0001x^2 \] \[C'(x) =2 - 0.0002x \]

OpenStudy (anonymous):

\[C'(970)=2-.0002\times 970=1.81\] rounded to the nearest penny